# Streaming Federated Learning with Markovian Data

### v1

@@ 5 Numerical Results
(b) $K=100$
(c) $K=1000$
@@
Figure 2: Gradient norm as a function of the number of communication rounds for Local SGD, Minibatch SGD & Local SGD-M, with $\gamma=0.01,\eta=0.001$, $\beta=0.1$, $\lambda=0.01$, $\tau_{m}=100$ for all $m$ and $K=10$.
In this section, we experimentally evaluate the performance of Minibatch SGD, local SGD, and local SGD-M on a non-convex linear regression problem with Markovian data. In the following experiments, we study the average performance over $10$ random seeds in a setting similar to Dorfman and Levy [2022].
Let $\mathcal{U}\left(a,b\right)$ denote the uniform probability distribution over $[a,b]$. The data stream for each client $m$ is generated by a two-state Markov chain $(i_{t}^{(m,k)},\leavevmode\nobreak\ k\in[K],t\in\mathbb{N})$, where the probability of jumping from one state to another is $p\in\left(0,1\right)$, with mixing time $\Theta(1/p)$. [^2] Here we use the standard Big-$\Theta$ notation. Associated with each state $i\in\{0,1\}$ are the vectors $V_{m,i}\in\mathbb{R}^{10}$ which are drawn randomly for each seed according to $\mathcal{U}\left(0,1\right)$. For each seed and $i\in\{0,1\}$, half of the optimal parameters $w_{m,i}\in\mathbb{R}^{10}$ take the value $w_{i}^{1}\sim\mathcal{U}(0,1)$, while the other half take the value $w_{i}^{2}\sim\mathcal{U}(1,2)$.
The samples associated with client $m$ correspond to the observations
$x_{t}^{(m,k)}=w_{m,i_{t}^{(m,k)}}^{T}V_{m,i_{t}^{(m,k)}}+\epsilon_{t}^{(m,k)},$ (34)
where $\epsilon_{t}^{(m,k)}\sim\mathcal{U}(0,0.01),\leavevmode\nobreak\ t\in\mathbb{N},m\in[M],k\in[K]$. Since the Markov chain $(i_{t}^{(m,k)},\leavevmode\nobreak\ k\in[K],t\in\mathbb{N})$ is symmetric, the stationary distribution for every client $m$ is uniform; i.e., $\pi_{m}(0)=\pi_{m}(1)=1/2$. The local objective function for each client $m$ is then given by
$F_{m}(w)=$ $\frac{1}{2}\left[\left(w-w_{m,1}\right)^{T}V_{m,1}\right]^{2}$ (35)
where $r(w)=\frac{1}{2}\sum_{d=1}^{10}\frac{w_{[d]}^{2}}{1+w_{[d]}^{2}}$ is a non-convex regularizer often used in robust non-convex and smooth stochastic optimization [Tran et al., 2021, Fatkhullin et al., 2023].
Impact of Heterogeneity: In Figure 1, we plot the trajectories of the gradient norm $\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}$ over the communications rounds for each of the three algorithms, with different number of samples per communication round. As predicted by our analysis, the performance of Local SGD stalls when the number of local updates increases, while Minibatch SGD and Local SGD-M are more robust to the heterogeneity with improved performance as the number of samples $K$ in each communication round is increased.
Impact of Number of Clients $M$: In Figure 2, we plot the trajectories of the gradient norm $\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}$ for each of the three algorithms, with different numbers of clients. The gradient norm for all three algorithms decreases when the number of clients increases, indicating that collaboration is beneficial even with heterogeneous Markovian data.
@@ 6 Conclusion
Statistical dependence is a reality in data streams arising from physical and biological systems. A key question is whether collaboration remains beneficial for FL with Markovian data streams. To address this question, we showed that there is a speed-up in FL with smooth non-convex objectives for Minibatch SGD, Local SGD, and Local SGD with Momentum. However, there is a cost associated with Markovian data streams, quantified by an increase in the sample complexity compared with i.i.d. sampling.
@@ A.1 Pseudo-code for Minibatch SGD and Local SGD
        Sample $(x_{t}^{m,k})_{k=1}^{K}$
        for $k=0$ to $K-1$ do
           Compute $g_{t}^{(m,k)}=\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)$
@@
        Initialize the local model $w_{t}^{(m,0)}=w_{t}$
        Sample $(x_{t}^{m,k})_{k=1}^{K}$
        for $k=0$ to $K-1$ do
           $w_{t}^{(m,k+1)}=w_{t}^{(m,k)}-\eta\nabla f_{m}\left(w_{t}^{(m,k)};x_{t}^{(m,k)}\right)$
@@
     Server update: $w_{t+1}=\frac{1}{M}\sum_{m=1}^{M}w_{t}^{(m,K)}$
@@
A.2 Useful notation
The following notation will be utilized throughout the appendix.
@@ A.2 Useful notation
$s_{t}=\mathbb{E}\left[\left\lVert w_{t+1}-w_{t}\right\rVert^{2}\right].$ (39)
We use $\sum_{m,k}$ for $\sum_{m=1}^{M}\sum_{k=1}^{K}$ unless explicitly stated differently.
With this notation, the update rule of Minibatch SGD is
@@
the update rule of Local SGD is
@@
where $\tilde{\eta}_{t}=K\eta_{t}$, and the update rule of Local SGD-M is
@@ Appendix B Markov chain Preliminaries
$\left\lVert\mu-\nu\right\rVert_{TV}=2\sup_{A\in\mathcal{B}}\left|\mu(A)-\nu(A)\right|=\int_{\Omega}\left|d\mu-d\nu\right|.$ (43)
If $\mu$ and $\nu$ have well defined densities $g,h$ with respect to an underlying measure $\rho$, then:
@@
Below we give some equivalent definitions of the total variation distance:
@@
  • (i)
    We have:
[…]
  • (ii)
    Let $Y,Z$ be a coupling of $\mu$ and $\nu$, i.e. the marginal distribution of $Y$ is $\mu$ and the marginal distribution of $Z$ is $\nu$. Then:
    $\left\lVert\mu-\nu\right\rVert_{TV}=2\inf\left\{P\left(Y\neq Z\right),\text{$\left(Y,Z\right)$ is a coupling of $\mu,\nu$}\right\}.$ (46)
    The coupling $(Y,Z)$ that satisfies $\left\lVert\mu-\nu\right\rVert_{TV}=P\left(Y\neq Z\right)$ is then called the maximal coupling of $\mu$ and $\nu$.
The proof can be found in Proposition 3, [Roberts and Rosenthal, 2004].
Now we present some nice properties on the total variation distance w.r.t the stationary distribution of Markov chains:
Proposition B.3.
Let $X$ be a Markov chain on $\left(\Omega,\mathcal{B}\right)$ with transition kernel $P$ and stationary distribution $\pi$, then:
@@
    Let $d(n)=\sup_{x\in X}\left\lVert P^{n}\left(x,.\right)-\pi(.)\right\rVert_{TV}$. Then for $\mu$ varying over probability distribution defined over $\left(\Omega,\mathcal{B}\right)$:
    $d(n)=\sup_{\mu}\left\lVert\mu P^{n}-\pi\right\rVert_{TV}.$ (49)
@@
    Let $\left\{X_{0},X_{1},\dots,X_{n}\right\}$ be a finite sequence from the chain. Let $\mu,\nu$ two probability distribution defined over $\left(\Omega,\mathcal{B}\right)$, and $\mathcal{L}_{\mu},\mathcal{L}_{\nu}$ be the joint distribution of $\left\{X_{1},X_{2},\dots,X_{n}\right\}$ with $X_{1}\sim\mu$ and $X_{1}\sim\nu$ respectively. Then we have:
    $\left\lVert\mathcal{L}_{\mu}-\mathcal{L}_{\nu}\right\rVert_{TV}=\left\lVert\mu-\nu\right\rVert_{TV}.$ (51)
  • (v)
    For $i\in[n]$, let $\mu_{i}$ and $\nu_{i}$ be probability distributions on $\left(\Omega_{i},\mathcal{B}_{i}\right)$. Define $\mu=\bigotimes_{i=1}^{n}\mu_{i}$ and $\nu=\bigotimes_{i=1}^{n}\nu_{i}$ as two probability distributions on $(\Omega:=\bigtimes_{i=1}^{n}\Omega_{i},\mathcal{B}:=\bigotimes_{i=1}^{n}\mathcal{B}_{n})$. Then we have:
@@
The proof of $(i),(ii)$ and $(iii)$ can also be found in Proposition 3, [Roberts and Rosenthal, 2004]. Here we present the proof for the last two statements.
  • (iv)
    Consider the "cylinder set" $A=A_{0}\times A_{1}\times\dots A_{n}\in\bigotimes_{i=0}^{n}\mathcal{B}$. Then by the Markov property, we have:
    $\begin{split}\mathcal{L}_{\mu}\left(A\right)&=\int_{A_{0}}\int_{A_{1}}\dots\int_{A_{n-1}}\mu\left(dx_{0}\right)P(x_{0},dx_{1})\dots P\left(x_{n-1},A_{n}\right),\\ \mathcal{L}_{\nu}\left(A\right)&=\int_{A_{0}}\int_{A_{1}}\dots\int_{A_{n-1}}\nu\left(dx_{0}\right)P(x_{0},dx_{1})\dots P\left(x_{n-1},A_{n}\right).\end{split}$ (53)
    Hence,
    $\begin{split}&\left\lVert\mathcal{L}_{\mu}-\mathcal{L}_{\nu}\right\rVert_{TV}\\ &=2\sup_{A_{0}}\left|\int_{A_{0}}\mu\left(dx_{0}\right)-\nu\left(dx_{0}\right)\right|\sup_{A_{1},\dots A_{n}}\left|\int_{A_{1}}\dots\int_{A_{n-1}}P\left(x_{0},dx_{1}\right)\dots P\left(x_{n-1},A_{n}\right)\right|\\ &=\left\lVert\mu-\nu\right\rVert_{TV}\left|\int_{\Omega}\dots\int_{\Omega}P(x_{0},dx_{1})\dots P(x_{n-1},\Omega)\right|\\ &=\left\lVert\mu-\nu\right\rVert_{TV}.\end{split}$ (54)
  • (v)
@@
    where $\left(X_{m},Y_{m}\right)$ is the maximal coupling of $\mu_{m},\nu_{m}$. Let’s define $X=\left\{X_{m}\right\}_{m\in[M]}$ and $Y=\left\{Y_{m}\right\}_{m\in[M]}$. Then again by (46) we have
@@
    which follows from the union bound.
@@
Given a Markov chain defined on $\left(\Omega,\mathcal{B}\right)$, let $d(t)$ the total variation distance defined in (49), $\tau\left(\epsilon\right)$ and $\tau$ defined in Definition 3.3, for any positive integer $n$, we have
@@
Given $M$ independent time-homogeneous Markov chain, each defined on $\left(\Omega_{m},\mathcal{B}_{m}\right)$, with transition kernel $P_{m}$, stationary distribution $\pi_{m}$ and finite mixing time $\tau_{m}$, where $m\in[M]$. Consider the product Markov chain defined on $(\Omega:=\bigtimes_{m=1}^{M}\Omega_{m},\mathcal{B}:=\bigotimes_{m=1}^{M}\mathcal{B}_{m})$ with transition kernel given by the following tensor product:
$P\left(\left(x_{1},\dots,x_{M}\right),\left(A_{1},\dots,A_{M}\right)\right)=\prod_{m=1}^{M}P_{m}\left(x_{m},A_{m}\right)\quad\textrm{with $x_{m}\in\Omega_{m},A_{m}\in\mathcal{B}_{m}$, for all $m\in[M]$}.$ (60)
The stationary distribution of this product chain is $\pi=\bigotimes_{m=1}^{M}\pi_{m}$. Then the mixing time of this product Markov chain $\tau$ satisfies
@@
$\begin{split}d(t)&=\underset{x\in\mathcal{X}}{\sup}\left\lVert P^{t}\left(x,.\right)-\pi\right\rVert,\\ d_{m}(t)&=\underset{x\in\mathcal{X}_{m}}{\sup}\left\lVert P_{m}^{t}\left(x,.\right)-\pi_{m}\right\rVert.\end{split}$ (62)
@@
where $\tau\left(\epsilon\right):=\min\left\{t:d(t)\leq\epsilon\right\}$ and $\tau_{m}\left(\epsilon\right)=\min\left\{t:d_{m}(t)\leq\epsilon\right\}$. Hence, from (63),

### v2

@@ 5 Numerical Results
In this section, we evaluate the performance of Local SGD, Minibatch SGD, and Local SGD with Momentum on the Beijing Multi-Site Air-Quality dataset [7], which contains hourly measurements from 12 weather stations across China, collected between March 2013 and February 2017. The prediction target is seasonality-adjusted PM2.5 concentration, using other air quality and meteorological indicators as features. In Appendix E, we provide further details for the time series data preprocessing, non-convex objective, and additional experiments to investigate the impact of the Markov chain’s spectral gap with synthetic data.
We partition the data temporally, reserving the last 12 months for testing and using the preceding 36 months for training. To simulate a federated setting with a larger number of clients, we create virtual clients by randomly sampling contiguous windows of $n\in\{6,12\}$ months from the training period. We use a linear regression model with a non-convex regularizer to encourage sparsity, based on preliminary correlation analysis indicating that some features are not informative. The regularization suppresses less relevant features while preserving smoothness for optimization, and is widely used in robust non-convex optimization [57, 20].
@@
(b) $K=50$
(c) $K=100$
Figure 1: Gradient norm as a function of the number of communication rounds for Local SGD, Minibatch SGD, Local SGD-M, and SCAFFOLD, with $\gamma=0.1,\eta=0.01$, $\beta=0.5$, $\lambda=0.01$ for 120 clients (each client has access to 12 consecutive months of training data) and different numbers of local steps.
@@
Figure 2: Gradient norm as a function of the number of communication rounds for Local SGD, Minibatch SGD & Local SGD-M, with $\gamma=0.1,\eta=0.001$, $\beta=0.5$, $\lambda=0.01$ and $K=100$ for different numbers of clients. Each client has access to a window of 66 consecutive months of training data.
Impact of heterogeneity: In Figure 1, we plot the trajectories of the gradient norm over the communication rounds for Minibatch SGD, Local SGD, Local SGD-M, and SCAFFOLD [27], which is the first algorithm proposed in the i.i.d. setting to mitigate heterogeneity in FL. We compare these methods under varying numbers of samples, $K$, per communication round. We observe that the performance of Minibatch SGD and Local SGD-M consistently improves as $K$ increases, whereas Local SGD and SCAFFOLD exhibit little to no improvement. This is consistent with our theoretical findings, which identify heterogeneity as a limiting factor as $K$ increases for Local SGD, but not for Minibatch SGD or Local SGD-M. For SCAFFOLD, we argue that its advantage is clearer in the partial participation setting. [9] also shows that incorporating momentum into SCAFFOLD under the partial client participation setting further improves the convergence. Extending our current analysis to evaluate the combined effect of momentum and control variates in the partial participation setting with Markovian data is a promising direction for future research.
Impact of Number of Clients $M$: In Figure 2, we plot the trajectories of the gradient norm for each of the three algorithms with different numbers of clients. For all three methods, the gradient norm decreases as the number of clients increases, indicating that collaboration is beneficial even in the presence of heterogeneous Markovian data—a trend consistent with our theoretical analysis.
@@ 6 Conclusion
Statistical dependence is a reality in data streams arising from physical and biological systems. A key question is whether collaboration remains beneficial for FL with Markovian data streams. To address this question, we showed via analysis and experiments with pollution data that there is a speed-up in FL with smooth non-convex objectives for Minibatch SGD, Local SGD, and Local SGD with Momentum. However, there is a cost associated with Markovian data streams, quantified by an increase in the sample complexity compared with i.i.d. sampling.
Acknowledgments and Disclosure of Funding
This research was supported in part by the French government, through the 3IA Côte d’Azur Investments in the Future project managed by the National Research Agency (ANR) with the reference numbers ANR-19-P3IA-0002 and ANR-24-CE25-225, in part by the European Network of Excellence dAIEDGE under Grant Agreement Nr. 101120726, in part by the Groupe La Poste, sponsor of the Inria Foundation, in the framework of the FedMalin Inria Challenge, and in part by the EU HORIZON MSCA 2023 DN project FINALITY (G.A. 101168816). This work was also partially supported by the Inria-Nokia Challenge Learn-Net. Experiments were carried out using the Grid’5000 testbed supported by a scientific interest group hosted by Inria and including CNRS, RENATER and several Universities as well as other organizations.
Appendix
@@ A.1 Pseudo-code for Minibatch SGD and Local SGD
   for $k=0$ to K1K-1 do
    Compute:
    $g_{t}^{(m,k)}=\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)$
@@
   Initialize the local model:
   $w_{t}^{(m,0)}=w_{t}$
   for $k=0$ to K1K-1 do
    $w_{t}^{(m,k+1)}=w_{t}^{(m,k)}$
    $\eta\nabla f_{m}\left(w_{t}^{(m,k)};x_{t}^{(m,k)}\right)$
@@
  Server update:
  $w_{t+1}=\frac{1}{M}\sum_{m=1}^{M}w_{t}^{(m,K)}$
@@
A.2 Notation
For clarity, we first recall some notations introduced in Section 3.
We use $M]=\left\{1,\dots,M\right\}$ to denote the set of positive integers up to $M$, and $K]_{0}=[0,K-1]$ to denote the set of non-negative integers less than $K$. For any vector $x\in\mathbb{R}^{d}$, we write $(x)_{i}$ for the $i$-th coordinate of $x$.
For every client $m\in[M]$, the data stream $X_{m}=(x_{t}^{m})_{t\in\mathbb{N}}$ is a time-homogeneous Markov process defined on state space $\Omega_{m}$ endowed with the corresponding Borel $\sigma$-field $\mathcal{B}_{m}$ and transition kernel. As an abuse of notation, for all $t\in\mathbb{N}$, and for all $k\in[K]_{0}$ we denote by $x_{t}^{(m,k)}\coloneqq x^{m}_{Kt+k}$ and by $x_{t}^{m}\coloneqq\left(x_{t}^{(m,k)}\right)_{k\in[K]_{0}}$, the $k$-th sample and the set of all $K$ samples used by client $m$ during round $t$, respectively.
The system-level Markov chain $X=(X_{m})_{m\in[M]}$ is a product Markov chain defined on $\left(\Omega\coloneqq\bigtimes_{m\in[M]}\Omega_{m},\mathcal{B}\coloneqq\bigotimes\mathcal{B}_{m}\right)$, which, at each time step, moves independently on each of $M$ coordinates according to the corresponding transition kernel. Its transition matrix is defined at the Kronecker tensor product $P=\bigotimes_{m\in[M]}P_{m}$. Similarly, to simplify the notation, we write $x_{t}^{k}:=\left(x_{t}^{(m,k)}\right)_{m\in[M]}$ and $x_{t}\coloneqq\left(x_{t}^{k}\right)_{k\in[K]_{0}}$.
For a Markov chain, we denote by $\mathbb{P}_{q}$ and $\mathbb{E}_{q}$ the corresponding probability distribution and expectation with initial distribution $q$. When $q=\delta(x)$, we simply write $\mathbb{P}_{x}$ and $\mathbb{E}_{x}$. We also denote by $\mathcal{F}_{t}\coloneqq\sigma\left(x_{i}^{(m,k)},m\in[M],k\in[K]_{0},i\in[t]_{0}\right)$ the $\sigma$-algebra generated by all data points received up to round $t$, and write $\mathbb{E}_{t}\!\left[.\right]$ as an alias for the conditional expectation $\mathbb{E}\!\left[.|\mathcal{F}_{t}\right]$.
The following notation will be used throughout the Appendix:
@@ A.2 Notation
$s_{t}=\mathbb{E}\left[\left\lVert w_{t+1}-w_{t}\right\rVert^{2}\right],$ (21)
where we use $\sum_{m,k}$ for $\sum_{m\in[M]}\sum_{k\in[K]_{0}}$ unless explicitly stated differently.
With this notation, the global update rule of Minibatch SGD is
@@
the global update rule of Local SGD is
@@
where $\tilde{\eta}_{t}=K\eta_{t}$, and the global update rule of Local SGD-M is
@@ Appendix B Markov Chain Preliminaries
In this section, we introduce some basic properties of Markov chains on general state spaces. In the following, we let $\Omega$ be a Polish space, and $\mathcal{B}$ denote the associated Borel $\sigma$-field.
We begin with the following formal definition of transition kernels.
Definition B.1.
If $P=\left\{P(x,A),x\in\Omega,A\in\mathcal{B}\right\}$ such that:
  • (i)
    for each $A\in\mathcal{B}$, $P\left(.,A\right)$ is a non-negative measurable function on $\Omega$,
  • (ii)
    for each $x\in\Omega$, $P\left(x,.\right)$ is a probability measure on $\mathcal{B}$,
then $P$ is called a transition kernel.
If $\Omega$ is countable, then P(x,y)P(x,{y}) corresponds to the transition probability of moving from state $x$ to state $y$ in a single step. However, in the case of continuous state spaces, we may have $P\left(x,{y}\right)=0$ for all $y\in\Omega$, hence the need of defining $P\left(x,A\right)$ as the probability of jumping from the current state $x$ into a subset $A\subseteq\Omega$.
Now we are ready to define a Markov chain.
Definition B.2.
A time-homogeneous Markov chain defined on $\left(\Omega,\mathcal{B}\right)$ with transition kernel $P$ is a sequence of random variables $\left(X_{i}\right)_{i\in\mathbb{N}}$ taking values in $\Omega$ such that for all $i\in\mathbb{N}$:
$\mathbb{P}\left(X_{i}\in A\mid X_{i-1}=x_{i-1};X_{i-2}=x_{i-2};\dots;X_{1}=x_{1}\right)$ $=\mathbb{P}\left(X_{i}\in A\mid X_{i-1}=x_{i-1}\right)$ (25)
$=P(x_{i-1},A).$ (26)
The Markov property (25) can be expressed in the following equivalent form:
Proposition B.3 (39, Proposition 3.4.3).
Given a time-homogeneous Markov chain $(X_{i})_{i\in\mathbb{N}}$ on $\left(\Omega,\mathcal{B}\right)$, and $f:\Omega\mapsto\mathbb{R}$ a bounded and measurable function, then:
$\mathbb{E}\!\left[f\left(X_{n+1},X_{n+2},\dots\right)\mid X_{n}=x_{n};X_{n-1}=x_{n-1};\dots;X_{1}=x_{1}\right]$ (27)
$=\mathbb{E}_{x_{n}}\!\left[f\left(X_{1},X_{2},\dots\right)\right]$ (28)
We recall the definition of a stationary measure as follows:
Definition B.4.
The stationary measure of a time-homogeneous Markov chain defined on $(\Omega,\mathcal{B})$ with transition kernel $P$ is a probability measure $\pi$ such that, for all $A\in\mathcal{B}$:
$\pi(A)=\int_{\Omega}\mathrm{d}\pi(x)P(x,A),$ (29)
or, equivalently:
$\mathrm{d}\pi(y)=\int_{\Omega}\mathrm{d}\pi(x)P(x,\mathrm{d}y).$ (30)
Before formally defining the system-level Markov chain introduced in Section 3.2, we need the following result on the measurability of the product of measurable functions.
Lemma B.5.
Let $\left(\Omega_{1},\mathcal{B}_{1}\right)$ and $\left(\Omega_{2},\mathcal{B}_{2}\right)$ be two measurable spaces. Let $f$ and $g$ be two real-valued measurable functions on $\Omega_{1}$ and $\Omega_{2}$, respectively. Let $h$ be a real-valued function defined as:
$h:\Omega_{1}\times\Omega_{2}$ $\rightarrow\mathbb{R}$ (31)
$\left(x_{1},x_{2}\right)$ $\mapsto h(x_{1},x_{2})=f(x_{1})g(x_{2}).$ (32)
Then $h$ is $\mathcal{B}_{1}\otimes\mathcal{B}_{2}$-measurable.
Proof.
Consider $\tilde{f}$ and $\tilde{g}$ defined as:
$\tilde{f}:\Omega_{1}\times\Omega_{2}$ $\rightarrow\mathbb{R}$ (33)
$(x_{1},x_{2})$ $\mapsto\tilde{f}(x_{1},x_{2})=f(x_{1})$ (34)
and
$\tilde{g}:\Omega_{1}\times\Omega_{2}$ $\rightarrow\mathbb{R}$ (35)
$(x_{1},x_{2})$ $\mapsto\tilde{g}(x_{1},x_{2})=g(x_{2}).$ (36)
As $f$ and $g$ are $\mathcal{B}_{1}$-measurable and $\mathcal{B}_{2}$-measurable, $\tilde{f}$ and $\tilde{g}$ are $\mathcal{B}_{1}\otimes\mathcal{B}_{2}$-measurable. Indeed,
$\tilde{f}^{-1}\left([a,+\infty)\right)$ $=f^{-1}\left([a,+\infty)\right)\times\Omega_{2}\in\mathcal{B}_{1}\otimes\mathcal{B}_{2},~\text{and}$ (37)
$\tilde{g}^{-1}\left([a,+\infty)\right)$ $=\Omega_{1}\times g^{-1}\left([a,+\infty)\right)\in\mathcal{B}_{1}\otimes\mathcal{B}_{2}.$ (38)
As $h=\tilde{f}\tilde{g}$, and the product of two measurable functions on the same measurable space is measurable [10, Proposition 2.1.7], we proved that $h$ is $\mathcal{B}_{1}\otimes\mathcal{B}_{2}$-measurable. ∎
The above result can be generalized to the product of more than two functions in a trivial way. We omit the proof for readability.
Now, we provide a formal definition of the system-level Markov chain introduced in Section 3.2. We show that it is indeed a well-defined Markov chain and characterize its transition kernel and stationary distribution.
Proposition B.6.
Consider $M$ time-homogeneous Markov chains $X_{m},m\in[M]$, each defined on $\left(\Omega_{m},\mathcal{B}_{m}\right)$ with transition kernel $P_{m}$ and stationary distribution $\pi_{m}$, respectively. The product Markov chain $X$, evolving on $\left(\Omega=\bigtimes_{m\in[M]}\Omega_{m},\mathcal{B}=\bigotimes_{m\in[M]}\mathcal{B}_{m}\right)$, with transition kernel $P$ defined as:
$P\left(x,A\right)=\prod_{m\in[M]}P_{m}(x_{m},A_{m}),$ (39)
$\text{for all }x=(x_{m})_{m\in[M]}\in\Omega~\text{and }A=\left(\bigtimes_{m\in[M]}A_{m}\right)\in\mathcal{B},$ (40)
is a well-defined Markov chain with stationary distribution $\pi=\bigotimes_{m\in[M]}\pi_{m}$.
Proof.
By construction, it is easy to see that for each $x\in\Omega$, $P(x,.)$ is a product measure on $\mathcal{B}$, hence the second condition in Definition B.4 is satisfied. Also by construction, for every $A\in\mathcal{B}$, $P(.,A)$ is non-negative, and Lemma B.5 shows that $P(.,A)$ is $\mathcal{B}$-measurable. Thus, $P$ also satisfies the first condition and is therefore a well-defined transition kernel.
We still need to show that the Markov property (25) is achieved. For simplicity, we will do so for the case $M=2$, but extension to any $M<\infty$ is straightforward.
Denote by $X_{1}=\left(X^{1}_{i}\right)_{i\in\mathbb{N}}$ and $X_{2}=\left(X^{2}_{i}\right)_{i\in\mathbb{N}}$ the two chains. Indeed, as the chains are independent, we have that:
$\mathbb{P}\left[\left(X^{1}_{i},X^{2}_{i}\right)\in A_{1}\times A_{2}\;\middle|\;\left(X^{1}_{i-1},X^{2}_{i-1}\right)=(x^{1}_{i-1},x^{2}_{i-1}),\dots,\left(X^{1}_{1},X^{2}_{1}\right)=\left(x^{1}_{1},x^{2}_{1}\right)\right]$ (41)
$=\mathbb{P}\left(X^{1}_{i}\in A_{1}\;\middle|\;X^{1}_{i-1}=x^{1}_{i-1},\dots,X^{1}_{1}=x^{1}_{1}\right)~\mathbb{P}\left(X^{2}_{i}\in A_{2}\;\middle|\;X^{2}_{i-1}=x^{2}_{i-1},\dots,X^{2}_{1}=x^{2}_{1}\right)$ (42)
$=P_{1}\left(x^{1}_{i-1},A_{1}\right)~P_{2}\left(x^{2}_{i-1},A_{2}\right)$ (43)
$=P\left((x^{1}_{i-1},x^{2}_{i-1}),(A_{1}\times A_{2})\right).$ (44)
Since for every $A\in\mathcal{B}$, $P(.,A)$ is non-negative and $\mathcal{B}$-measurable, a direct application of Tonelli’s Theorem gives us the statement on the stationary measure of the product chain:
$\pi(A)$ $=\left(\bigotimes_{m\in[M]}\pi_{m}\right)(A)=\prod_{m\in[M]}\pi_{m}\left(A_{m}\right)$ (45)
$=\int_{\Omega_{1}}\mathrm{d}\pi_{1}(x_{1})P_{1}(x_{1},A_{1})\dots\int_{\Omega_{M}}\mathrm{d}\pi_{M}(x_{M})P_{M}(x_{M},A_{M})$ (46)
$=\int_{\Omega_{1}}\dots\int_{\Omega_{m}}P_{1}\left(x_{1},A_{1}\right)\dots P_{M}\left(x_{M},A_{M}\right)\mathrm{d}\pi_{1}(x_{1})\dots\mathrm{d}\pi_{M}(x_{M})$ (47)
$=\int_{\Omega}\mathrm{d}\pi\left(x\right)P\left(x,A\right).$ (48)
@@
$\left\lVert\mu-\nu\right\rVert_{TV}=\sup_{A\in\mathcal{B}}\left|\mu(A)-\nu(A)\right|=\frac{1}{2}\int_{\Omega}\left|\mathrm{d}\mu-\mathrm{d}\nu\right|.$ (49)
Below, we give an equivalent definition of the total variation distance:
Proposition B.8.
Let $\left(Y,Z\right)$ be a coupling of $\mu$ and $\nu$, i.e. the marginal distribution of $Y$ is $\mu$ and the marginal distribution of $Z$ is $\nu$. Then:
@@
The coupling $(Y,Z)$ that satisfies $\left\lVert\mu-\nu\right\rVert_{TV}=\mathbb{P}\left(Y\neq Z\right)$ is then called the maximal coupling of $\mu$ and $\nu$.
The proof can be found in [48, Proposition 3].
Now we present some nice properties on the total variation distance with respect to the stationary measure of Markov chains:
@@
Let $X_{1},X_{2},\dots$ be a Markov chain on $\left(\Omega,\mathcal{B}\right)$ with transition kernel $P$ and stationary distribution $\pi$, then:
@@
    Let $d(n)=2\sup_{x\in\Omega}\left\lVert P^{n}\left(x,.\right)-\pi(.)\right\rVert_{TV}$. Then for $\mu$ varying over probability distribution defined over $\left(\Omega,\mathcal{B}\right)$:
    $d(n)=2\sup_{\mu}\left\lVert\mu P^{n}-\pi\right\rVert_{TV}.$ (53)
@@
    For $i\in[n]$, let $\mu_{i}$ and $\nu_{i}$ be probability distributions on $\left(\Omega_{i},\mathcal{B}_{i}\right)$. Define $\mu\coloneqq\bigotimes_{i=1}^{n}\mu_{i}$ and $\nu\coloneqq\bigotimes_{i=1}^{n}\nu_{i}$ as two probability distributions on $(\Omega\coloneqq\bigtimes_{i=1}^{n}\Omega_{i},\mathcal{B}\coloneqq\bigotimes_{i=1}^{n}\mathcal{B}_{n})$. Then we have:
@@
The proof of $(i),(ii)$ and $(iii)$ can also be found in [48, Proposition 3]. Here we present the proof for the last statements.
@@
where, for all $m\in[M]$, $\left(X_{m},Y_{m}\right)$ is the maximal coupling of $\mu_{m},\nu_{m}$. Define $X=\left\{X_{m}\right\}_{m\in[M]}$ and $Y=\left\{Y_{m}\right\}_{m\in[M]}$. Then again by (50) we have
@@
which follows from the union bound. ∎
We now give the definition of uniform geometric ergodicity, which is commonly used in the literature of stochastic optimization with Markovian data [3, 37].
Definition B.10.
A Markov chain defined on $\left(\Omega,\mathcal{B}\right)$ with transition kernel $P$ is uniformly geometrically ergodic if:
$\sup_{x\in\Omega}\left\lVert P^{n}\left(x,.\right)-\pi\right\rVert_{\mathrm{TV}}\leq c\rho^{n},$ (58)
for some $c\leq\infty$ and $\rho<1$.
Uniform geometric ergodicity implies that, regardless of the starting distribution, the chain converges exponentially fast to its stationary measure in terms of total variation norm.
To end this section, we have the following results on the mixing time of Markov chains.
@@
Given a Markov chain defined on $\left(\Omega,\mathcal{B}\right)$, let $d(t)$ the total variation distance defined in (53), the mixing time $\tau\left(\epsilon\right)$ and $\tau$ are defined as:
$\tau(\epsilon)$ $=\min\{t:d(t)\leq\epsilon\},$ (59)
$\tau$ $=\tau(1/4),$ (60)
For any positive integer $n$, we have
@@
Consider the product Markov chain in Proposition B.6. The mixing time of this product Markov chain $\tau$ (see definitions in Proposition B.11) satisfies
@@
$\begin{split}d(t)&=2\underset{x\in\Omega}{\sup}\left\lVert P^{t}\left(x,.\right)-\pi\right\rVert,\\ d_{m}(t)&=2\underset{x\in\Omega_{m}}{\sup}\left\lVert P_{m}^{t}\left(x,.\right)-\pi_{m}\right\rVert.\end{split}$ (65)
@@
where $\tau\left(\epsilon\right)\coloneqq\min\left\{t:d(t)\leq\epsilon\right\}$ and $\tau_{m}\left(\epsilon\right)=\min\left\{t:d_{m}(t)\leq\epsilon\right\}$. Hence, from (66),